OpenStudy (anonymous):

will give medals The volume of a rectangular prism can be computed by using the formula V = lwh. Which shows an equation solving for w? A. w=V/lh B. w=lh/V C. w=Vl/h D. w=Vlh

OpenStudy (anonymous):

its just like an alegbra expression solving for a number. you have to isolate W

OpenStudy (anonymous):

?

OpenStudy (anonymous):

put it by itself

OpenStudy (anonymous):

you would have to divide l and h from V to put W by itsself

OpenStudy (anonymous):

Which is basically your answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ello love :D need some help?

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

sorry i couldnt help as much as i wish

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

okay so we need to find w so we isolate it by changing the original equation: V=lwh correct?

OpenStudy (anonymous):

no its w=....

OpenStudy (anonymous):

duh! haha but the original equation is V=lwh so we simply divide by l and h

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

yea! :D its that easy haha

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

okay so when your isolating (solving) for a specific variable, you must isolate it meaning you subtract everything added, add everynegative or subtraction till you isolate the variable.... do you want practice? illl post problems and walk you through it :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay :) so we'll start with a simply one: 8+x=10 solve for x and show me the steps.... can you do it?

OpenStudy (anonymous):

2=x

OpenStudy (anonymous):

yes common sense i know but yo solve for this showing steps we subtract 8 from both sides which leaves us with x=2 8+x=10 8-8+x=10-8 x=2

OpenStudy (anonymous):

oh okay but idk about my question

OpenStudy (anonymous):

well its the same thing, simply do the opposite of the thing you want to move: V=lwh V/lh=lwh/lh V/lh=w

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no problem :)